Greenergy has the opportunity to license a technology, WiCharge?, that through a variety of “green” sources (including solar, body heat, motion of the user, etc.) allows handheld devices such as tablets and cell phones to function without being plugged in for 400% longer than the battery alone will allow. The prototypes the firm has developed are designed to fit as a cover, where the cover protects and also “plugs” into the power input for the device.

The demand for such a product is expected to be explosive, as no competing device will be ready for the market for at least 2 years. First year sales figures estimate that 1 million devices can be sold for an average of $80 each. Sales volume will grow by 75% in the second year, but then will decrease by 50% in year 3 and no devices will be sold in the fourth year. The devices will be sold at $80 in all 3 years.

The margins on the initial sale are very high (around 70%), but one drawback is that the EPA has mandated that Greenergy must collect and dispose of the devices because the materials (while safe for humans) have detrimental environmental effects if disposed of improperly. The devices will last 2 years, at which time the firm must pay an estimated $7.00 per device for disposal and $20 as an incentive to encourage customers to recycle the devices. Greenergy has agreed to pay the EPA a fine of $27 for every Greenergy device that is not returned by the end of year 5. They anticipate that 90% of devices will be returned for recycling very close to 2 years after the sale (those not returned by two years will never be returned), and they will pay the fine at the end of year 5 for the rogue devices.

The license fee to use the technology (which is owned by a prestigious university) is $77M upfront and $10M for every year devices are sold. Assume this license fees are immediately tax deductible at the time they are paid and the tax rate is 35% for Greenergy.

  1. What is the payback period? Would you decide to pay the license fee and undertake the project if your decision rule required payback within 3 years? 5 years?

  1. Use the excel IRR function with a guess of 3%, what is the IRR and what would your decision be if the appropriate discount rate is 10%? Repeat for a guess of 15%, what is the IRR and what would your decision be if the appropriate discount rate is 10%? Use the IRR investing rule only for your decision.

  1. According to the NPV decision rule, should the Greenergy managers agree to pay the licensing fees for WiCharge? if the appropriate discount rate is 10%? What if the appropriate discount rate is 15%?

  1. Use your answers for parts a, b and c to create a plot (x=Rate,y=NPV), what does the graph of the NPV for this project look like using discount rates from 0%-15%?

  1. Briefly change the total cash flow in year 0 to -$50.5M, what happens to your solution for IRR, and why?
Answer #1
A Sales in units in year 1              1,000,000         27,000,000
B=A*1.75 Sales in units in year 2              1,750,000         47,250,000
C=B*0.5 Sales in units in year 3                  875,000         23,625,000
D=A*0.9*(20+7) Cost of collection and disposal in year3 $24,300,000
E=B*0.9*27 Cost of collection and disposal in year4 $42,525,000
F=C*0.9*27 Cost of collection and disposal in year5 $21,262,500
N Year 0 1 2 3 4 5
G Unit Sales            1,000,000         1,750,000          875,000
H Sales price per unit $80 $80 $80
I=G*H Sales Revenue $80,000,000 $140,000,000 $70,000,000
J=0.7*I Annual margin(Befor tax) $56,000,000 $98,000,000 $49,000,000
K=J*(1-0.35) After tax annual Margin $36,400,000 $63,700,000 $31,850,000
L Cost of collection and disposal(Before tax) $24,300,000 $42,525,000 $21,262,500
M=L*(1-0.35) After tax cost of collection and disposal $15,795,000 $27,641,250 $13,820,625
P=0.1*(A+B+C)*27 Fine for devices not returned $        9,787,500
Q License Fees(Before tax) $77,000,000 $10,000,000 $10,000,000 $10,000,000
R=Q*(1-0.35) After tax license fees $50,050,000 $6,500,000 $6,500,000 $6,500,000
S=K-M-P-R NET CASH FLOW ($50,050,000) $29,900,000 $57,200,000 $9,555,000 ($27,641,250) ($23,608,125)
CUMULATIVE NET CASH FLOWS ($50,050,000) ($20,150,000) $37,050,000 $46,605,000 $18,963,750 ($4,644,375)
Cumulative cash flow is positive in 3 years
If required payback is within 3 years, the project should be undertaken
Cumulative cash flow is negative in 5 years
If required payback is within 5years, the project should NOT be undertaken
SUM
PV1=S/(1.03^N) Present Value of cash flow at 3% discount         (50,050,000)         29,029,126       53,916,486       8,744,179    (24,558,893)        (20,364,576) (3,283,678)
PV2=S/(1.1^N) Present Value of cash flow at 10% discount -50050000 27181818.18 47272727.27 7178812.923 -18879345.67 -14658788.21 (1,954,776)
PV3=S/(1.15^N) Present Value of cash flow at 15% discount -50050000 26000000 43251417.77 6282567.601 -15803974.4 -11737410.51 (2,057,400)
at 3% discount NPV=            (3,283,678)
at10% discount NPV=            (1,954,776)
at15% discount NPV=            (2,057,400)
In all cases NPV is negative
The project should not be undertaken